Glucose is transported into the intestinal cells via cotransport with sodium. Assume the
following conditions:
Temp = 37oC
Sodium (in) = 12.0 mM
Sodium (out) = 145 mM
Glucose (out) =28 μM
ΔΨ = -72mV (inside is negative)
a. What is the ΔG for sodium transport?
b. What is the upper limit of glucose transport?
Answer #1

ΔG (x) = RT ln (x2/x1) = 2.303 RT log10 (x2/x1)

where x is the concentartion

ΔG (x) = 2.303* 8.314 x 10−3 kJ mol−1 K−1 *310K*log10 ( 145/12 ) = 6.423 KJ. mol-1